# Monge solutions and uniqueness in multi-marginal optimal transport with hierarchical jumps

### v1

@@ Abstract
We introduce Hierarchical Jump multi-marginal transport (HJMOT), a generalization of multi-marginal optimal transport where mass can ”jump” over intermediate spaces via augmented isolated points. Established on Polish spaces, the framework guarantees the existence of Kantorovich solutions and, under sequential differentiability and a twist condition, the existence and uniqueness of Monge solutions. This core theory extends robustly to diverse settings, including smooth Riemannian manifolds, demonstrating its versatility as a unified framework for optimal transport across complex geometries.
@@ 1 Introduction
The theory of optimal transport (OT) fundamentally addresses the problem of efficiently moving mass between two spaces $X_{0}$ and $X_{K}$, minimizing a cost defined on pairs of points from these spaces (see [MR2459454]). The framework of multi-marginal optimal (MOT) theory extends this by introducing intermediate spaces $X_{1},\cdots,X_{K-1}$. It requires the mass to traverse a sequential path $X_{0}\to X_{1}\to\cdots\to X_{K-1}\to X_{K}$, minimizing a cumulative cost that is aggregated over each adjacent pair of spaces. However, this classical formulation imposes a rigid sequential structure, which may not be optimal in applications where direct jumps between non-adjacent spaces can lead to lower total cost. However, this classical formulation imposes a rigid sequential structure, which may not be optimal in applications where direct jumps between non-adjacent spaces can lead to lower total cost. To model this flexibility, we augment each intermediate space $X_{k}$ ($1\leq k\leq K-1$) with an isolated point $\partial_{k}$, defining an extended Polish space $\widehat{X}_{k}=X_{k}\cup{\partial_{k}}$. The origin space and terminal space remain unchanged, i.e. $\widehat{X}_{0}:=X_{0}$ and $\widehat{X}_{K}:=X_{K}$. A path traversing $\partial_{k}$ signifies that space $X_{k}$ is skipped in the transport leg between its neighbors. The active indices map $I(\omega)$ extracts the ordered sequence of spaces actually visited (excluding those skipped via $\partial_{k}$) from any path $\omega\in\Omega:=\prod_{k=0}^{K}\widehat{X}_{k}$. We leverage this structure in the following way: We define a path cost function $c(\omega)$, which has a form of sums of the pairwise costs between consecutive active indices identified by $I(\omega)$. The maximum adjacent cost $\tilde{c}_{i,i+1}$ provides an upper bound facilitating analysis. For fix $\mu_{0}\in\mathcal{P}(X_{0})$ and $\mu_{K}\in\mathcal{P}(X_{k})$, the core Kantorovich formulation seeks a coupling $\pi\in\Pi(\mu_{0},\cdots,\mu_{K})$ minimizing the expected path cost :
@@
The classical Monge problem with quadratic cost has been extensively studied in various geometric settings. In the Euclidean space, Brenier’s seminal theorem established that when the source measure is absolutely continuous with respect to the Lebesgue measure, there exists a unique optimal transport map for the squared distance cost [MR1100809]. This breakthrough was later extended to compact Riemannian manifolds by McCann [MR1844080], who proved existence and uniqueness of optimal maps under the same absolute continuity condition. Subsequently, Figalli [MR2459454] resolved the non-compact case, showing that the problem admits a solution on general complete non-compact Riemannian manifolds.
In Alexandrov spaces with lower curvature bounds, Bertrand [MR2442054] proved the existence and uniqueness of optimal maps for the quadratic cost under the assumption that the source measure is absolutely continuous with respect to the Hausdorff measure, provided the space is finite-dimensional and compact. Although this result represents an important milestone, the Monge problem in fully general Alexandrov spaces remains only partially understood. Beyond the Riemannian and Alexandrov settings, significant advances have been made in more general metric measure spaces. Gigli [MR2984123] obtained analogous results in synthetic Ricci curvature lower bound spaces (${\sf RCD}(K,N)$). Figalli and Rifford [MR2647137] resolved the existence and uniqueness of Monge solutions on sub-Riemannian manifolds for the squared distance cost under the assumption of absolute continuity of the source measure. The present work situates itself within this broad landscape. The proposed framework for the HJMOT problem not only encompasses the aforementioned results of Brenier, McCann, Figalli, and Bertrand, but also aligns with recent developments in general metric measure spaces. In this sense, our results provide a unified perspective that integrates and extends the classical theory of optimal transport across Euclidean, Riemannian, Alexandrov, and synthetic curvature settings.
@@
We topologize $\widehat{X}_{k}$ so that $\{\partial_{k}\}$ is isolated in $\widehat{X}_{k}$. Consequently, $\widehat{X}_{k}$ is a Polish space (see [MR1321597, Section 3]).
@@
Let $\{\pi_{n}\}$ be a minimizing sequence to (2.1) with respect to $c$. We say that cost function $c$ is coercive if there exists a measurable function $\Phi:\Omega\to[0,+\infty]$ such that
@@
    $\sup_{n}\int_{\Omega}\Phi\,d\pi_{n}<\infty.$
@@
Let $\{\pi_{n}\}\subset\mathcal{P}(\Omega)$ and $\pi_{n}\in\Pi(\mu_{0},\mu_{1}^{n},\cdots,\mu_{K-1}^{n},\mu_{K})$ be a minimizing sequence to (2.1) with $\mu_{k}^{n}={T_{k}}_{\#}\pi_{n}$ for $k=1,\cdots,K-1$. If cost function $c$ is coercive, then $\{\pi_{n}\}$ is uniformly tight.
@@
Take $\{\pi_{n}\}$ noted. By the coercivity condition, there exists a measurable function $\Phi:\Omega\to[0,+\infty]$ and a constant $C>0$ such that
@@
    $\sup_{n}\int_{\Omega}\Phi\,d\pi_{n}=C.$
@@
Let $\{\pi_{n}\}\subset\Pi(\mu_{0},\mu_{1}^{n},\cdots,\mu_{K-1}^{n},\mu_{K})$ be a minimizing sequence to (2.1) with $\mu_{k}^{n}={T_{k}}_{\#}\pi_{n}$ for $k=1,\cdots,K-1$. By definition of $M$, we have
@@ 3 Existence and Uniqueness for the Monge Problem
Let $\{\pi_{n}\}$ be a minimizing sequence to (2.1) with respect to $c$. We say that cost function $c$ is strongly coercive if there exists a measurable function $\Phi:\Omega\to[0,+\infty]$ such that
@@
    $\sup_{n}\int_{\Omega}\Phi\,d\pi_{n}<\infty.$
@@
    If there exists $\omega\in\Omega$ such that $\Psi(\omega)=\infty$, then $c(\omega)=\infty$.
In fact, strong coercivity is equivalent to: for any $R>0$, the set $\{\omega\in\Omega\mid c(\omega)\leq R\}$ is compact.
Lemma 3.1 (Kantorovich duality [MR4809472]).
@@
Since $c$ is non-negative semicountinuous and $\Pi(\mu_{0}^{*},\cdots,\mu_{K}^{*})$ is uniformly tight (Lemma 2.2), then by Kellerer’s theorem ([MR761565]) we have $D=M$ and there exists $v_{k}\in L^{1}(\widehat{X}_{k},\mu_{k}^{*})$ such that for evey $\omega\in\Omega$,
@@
Set $h(\widehat{x}_{0}):=\inf\{c(\omega)\mid\omega\in\Omega,\ T_{0}(\omega)=\widehat{x}_{0}\}$. If the cost function $c$ is strongly coercive, then $S(\widehat{x}_{0}):=\{\omega\in\Omega\mid T_{0}(\omega)=\widehat{x}_{0},\ c(\omega)=h(x_{0})\}$ is a non-empty compact set.
@@
For fix $\widehat{x}_{0}\in X_{0}$, by defined of $h(\widehat{x}_{0})$, there exists $\{\omega_{n}\}_{n=1}^{\infty}\subset\Omega$ such that
@@
Then, for any $\varepsilon>0$, there exists $N\in\mathbb{N}$ such that for any $n>N$,
$\left|c(\omega_{n})-h(\widehat{x}_{0})\right|<\varepsilon.$
In particular, if we take $\varepsilon=1$, then there exists $N_{1}\in\mathbb{N}$ such that for any $n>N_{1}$, $\left|c(\omega_{n})-h(\widehat{x}_{0})\right|<1$. Since $c$ is strongly coercive, the set $K:=\{\omega\in\Omega\mid c(\omega)\leq h(\widehat{x}_{0})+1\}$ is compact. Since $\{\omega_{n}\}_{n\geq N_{1}}^{\infty}\subset K$ and $K$ is compact, there exists $\omega_{n_{k}}$ such that $\omega_{n_{k}}\to\omega^{*}\in K$ as $k\to\infty$. For each $k$, $T_{0}(\omega_{n_{k}})=\widehat{x}_{0}$ and $T_{0}$ is continuous map, then
@@
Lemma 3.4.
If $c$ is strongly coercive, then $h(\widehat{x}_{0})$ is lower semicontinuous on $X_{0}$.
Proof.
For any subsequences $\{\widehat{x}_{0}^{n}\}\subset X_{0}$ such that $\widehat{x}_{0}^{n}\to\widehat{x}_{0}$ as $n\to\infty$, if $\liminf_{n\to\infty}h(\widehat{x}_{0}^{n})=+\infty$, then $\liminf_{n\to\infty}h(\widehat{x}_{0}^{n})>h(\widehat{x}_{0})$. If $\liminf_{n\to\infty}h(\widehat{x}_{0}^{n})<\infty$, for each $n$, there exists $\omega_{n}\in\Omega$ such that
$T_{0}(\omega_{n})=\widehat{x}_{0}^{n}\quad\text{and}\quad c(\omega_{n})\leq h(\widehat{x}_{0}^{n})+\frac{1}{n}.$
Since $\liminf_{n\to\infty}h(\widehat{x}_{0}^{n})=L<+\infty$, there exists $R>0$ such that for a sufficiently large $n$, there is $h(\widehat{x}_{0}^{n})\leq R$. Therefore, $c(\omega_{n})\leq h(\widehat{x}_{0}^{n})+\frac{1}{n}\leq R+1$ for each sufficiently large $n$. By $c$ is strongly coercive, $\{\omega\in\Omega\mid c(\omega)\leq M+1\}$ is compact set and $\{\omega_{n}\}\subset K$ for each sufficiently large $n$, then there exists $\{\omega_{n_{k}}\}\subset\{\omega_{n}\}$ such that $\omega_{n_{k}}\to\omega^{*}\in\Omega$. Since $T_{0}$ is a continuous map, we have
$T_{0}(\omega)=\lim_{k\to\infty}T_{0}(\omega_{n_{k}})=\lim_{k\to\infty}\widehat{x}_{0}^{n_{k}}=\widehat{x}_{0}.$
By the difinition of $h$ and $c$ is lower semicontinuous,
[…]
Hence, $h$ is a semicontinuous function. ∎
@@
    The coupling $\pi^{*}=({\rm id},\,T^{*})_{\#}\mu_{0}$ is the unique solution of (2.1).
@@
exists in $\mathbb{R}$ and hence $c(\widehat{x}_{0}^{n},\widehat{x}_{1},\cdots,\widehat{x}_{K})\to c(\widehat{x}_{0},\widehat{x}_{1},\cdots\,\widehat{x}_{K})$ as $n\to\infty$. For any $\omega=(\widehat{x}_{0},\cdots,\widehat{x}_{K})\in S(\widehat{x}_{0})$, $c(\widehat{x}_{0}^{n},\widehat{x}_{1},\cdots,\widehat{x}_{K})-c(\widehat{x}_{0},\widehat{x}_{1},\cdots\,\widehat{x}_{K})\geq h(\widehat{x}_{0}^{n})-h(\widehat{x}_{0})$. Then, we have $c(\widehat{x}_{0}^{n},\widehat{x}_{1},\cdots,\widehat{x}_{K})-c(\widehat{x}_{0},\widehat{x}_{1},\cdots\,\widehat{x}_{K})+h(\widehat{x}_{0})\geq h(\widehat{x}_{0}^{n})$. Take the upper limit of both sides, $\limsup_{n\to\infty}h(\widehat{x}_{0}^{n})\leq h(\widehat{x}_{0})$. Since $h$ is lower semicontinuous, $\lim_{n\to\infty}h(\widehat{x}_{0}^{n})=h(\widehat{x}_{0})$. According to the definition of $h$, we can know that
$\frac{c(\widehat{x}_{0}^{n},\widehat{x}_{1},\cdots,\widehat{x}_{K})-c(\widehat{x}_{0},\widehat{x}_{1},\cdots\,\widehat{x}_{K})}{{\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0})}\geq\frac{h(\widehat{x}_{0}^{n})-h(\widehat{x}_{0})}{{\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0})}.$
On the other hand, due to
$\frac{c(\widehat{x}_{0}^{n},\widehat{x}_{1},\cdots,\widehat{x}_{K})-c(\widehat{x}_{0},\widehat{x}_{1},\cdots\,\widehat{x}_{K})}{{\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0})}\leq\frac{h(\widehat{x}_{0}^{n})+o({\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0}))-h(\widehat{x}_{0})}{{\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0})}.$
@@
and $D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}c(\omega)=D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}h$ for any $\omega\in S(\widehat{x}_{0})$. Suppose there exists $\omega_{1}\in S(\widehat{x}_{0})\cap{\rm supp}(\pi^{*})$ and $\omega_{2}\in S(\widehat{x}_{0})\cap{\rm supp}(\pi^{*})$ such that $\omega_{1}\neq\omega_{2}$. Since $\omega_{1}$ and $\omega_{2}$ are not equal, by $c$ is sequential twisted, $D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}c(\omega_{1})\neq D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}c(\omega_{2})$. But $\omega_{1},\omega_{2}\in S(\widehat{x}_{0})\cap{\rm supp}(\pi^{*})$,
@@
contradiction. Then $S(\widehat{x}_{0})\cap{\rm supp}(\pi^{*})$ is a singleton. In particular, if Polish spaces $X_{0}$ that lack a nontrivial curve structure, $S(\widehat{x}_{0})=\mathcal{F}(\widehat{x}_{0})$ is also a singleton. For each $\widehat{x}_{0}\in X_{0}$, define $T^{*}(\widehat{x}_{0})$ as the unique element of $S(\widehat{x}_{0})\cap{\rm supp}(\pi^{*})$. Consider
$G_{r}(T^{*}):$ $=\{(\widehat{x}_{0},\omega)\in X_{0}\times\Omega\mid\omega\in S(\widehat{x}_{0})\cap{\rm supp}(\pi^{*})\}$
@@
Since $\{(\widehat{x}_{0},\omega)\in X_{0}\times\Omega\mid\omega\in{\rm supp}(\pi^{*}),\,T_{0}(\omega)=\widehat{x}_{0}\}$ is a closed set and $\{(\widehat{x}_{0},\omega)\in X_{0}\times\Omega\mid c(\omega)=h(\widehat{x}_{0})\}$ is a Borel set, then $G_{r}$ is a Borel set and hence $T^{*}$ is a Borel map (see [MR188994]). By the disintegration theorem, for any Borel set $A\subset\Omega$
@@
where $\delta_{x_{0}}$ is a Dirac measure and $\kappa_{\widehat{x}_{0}}$ is the conditional probability measure of $\omega$ given $\widehat{x}_{0}$. Since $\kappa_{\widehat{x}_{0}}(S(\widehat{x}_{0})\cap\text{supp}(\pi^{*}))=1$ and $S(\widehat{x}_{0})\cap{\rm supp}(\pi^{*})$ is a singleton, then $\kappa_{\widehat{x}_{0}}=\delta_{T^{*}(\widehat{x}_{0})}$. Therefore
$\pi^{*}(A)=\int_{X_{0}}[\delta_{\widehat{x}_{0}}\otimes\kappa_{\widehat{x}_{0}}](A)\ \mu_{0}({\rm d}\widehat{x}_{0})=({\rm id},\,T^{*})_{\#}\mu_{0}(A).$
@@
Similarly, $S(\widehat{x}_{0})\cap{\rm supp}(\pi^{\prime})$ is also a singleton and $\kappa_{\widehat{x}_{0}}(S(\widehat{x}_{0})\cap\text{supp}(\pi^{*}))=1$, then $\kappa^{\prime}_{\widehat{x}_{0}}=\delta_{T^{*}(\widehat{x}_{0})}=\kappa_{\widehat{x}_{0}}$. Hence,

### v2

@@ Abstract
We introduce Hierarchical Jump multi-marginal transport (HJMOT), a generalization of multi-marginal optimal transport where mass can “jump” over intermediate spaces via augmented isolated points. Established on Polish spaces, the framework guarantees the existence of Kantorovich solutions and, under sequential differentiability and a twist condition, the existence and uniqueness of Monge solutions. This core theory extends robustly to diverse settings, including smooth Riemannian manifolds, demonstrating its versatility as a unified framework for optimal transport across complex geometries.
@@ 1 Introduction
The theory of optimal transport (OT) fundamentally addresses the problem of efficiently moving mass between two spaces $X_{0}$ and $X_{K}$, minimizing a cost defined on pairs of points from these spaces (see [16]). The framework of multi-marginal optimal (MOT) theory extends this by introducing intermediate spaces $X_{1},\cdots,X_{K-1}$. It requires the mass to traverse a sequential path $X_{0}\to X_{1}\to\cdots\to X_{K-1}\to X_{K}$, minimizing a cumulative cost that is aggregated over each adjacent pair of spaces. However, this classical formulation imposes a rigid sequential structure, which may not be optimal in applications where direct jumps between non-adjacent spaces can lead to lower total cost. To model this flexibility, we augment each intermediate space $X_{k}$ ($1\leq k\leq K-1$) with an isolated point $\partial_{k}$, defining an extended Polish space $\widehat{X}_{k}=X_{k}\cup{\partial_{k}}$. The origin space and terminal space remain unchanged, i.e. $\widehat{X}_{0}:=X_{0}$ and $\widehat{X}_{K}:=X_{K}$. A path traversing $\partial_{k}$ signifies that space $X_{k}$ is skipped in the transport leg between its neighbors. The active indices map $I(\omega)$ extracts the ordered sequence of spaces actually visited (excluding those skipped via $\partial_{k}$) from any path $\omega\in\Omega:=\prod_{k=0}^{K}\widehat{X}_{k}$. We leverage this structure in the following way: We define a path cost function $c(\omega)$, which has a form of sums of the pairwise costs between consecutive active indices identified by $I(\omega)$. The maximum adjacent cost $\tilde{c}_{i,i+1}$ provides an upper bound facilitating analysis. For fix $\mu_{0}\in\mathcal{P}(X_{0})$ and $\mu_{K}\in\mathcal{P}(X_{k})$, the core Kantorovich formulation seeks a coupling $\pi\in\Pi(\mu_{0},\cdots,\mu_{K})$ minimizing the expected path cost :
@@
The framework is equally applicable to geometric spaces lacking a global orientation. In Chapter 5, HJMOT is implemented on compact nonorientable Riemannian manifolds. By incorporating the orientable double cover as an intermediate space and allowing mass to jump selectively through this cover or to be transported directly, the Monge problem on a nonorientable manifold is reduced to an orientable setting. This concretely illustrates the modeling flexibility of HJMOT: by appropriately introducing an extended intermediate space adapted to the structure of the problem, the framework can decompose a geometrically intricate transport problem into more tractable subproblems.
The classical Monge problem with quadratic cost has been extensively studied in various geometric settings. In the Euclidean space, Brenier’s seminal theorem established that when the source measure is absolutely continuous with respect to the Lebesgue measure, there exists a unique optimal transport map for the squared distance cost [4]. This breakthrough was later extended to compact Riemannian manifolds by McCann [13], who proved existence and uniqueness of optimal maps under the same absolute continuity condition. Subsequently, Figalli [16] resolved the non-compact case, showing that the problem admits a solution on general complete non-compact Riemannian manifolds.
In Alexandrov spaces with lower curvature bounds, Bertrand [3] proved the existence and uniqueness of optimal maps for the quadratic cost under the assumption that the source measure is absolutely continuous with respect to the Hausdorff measure, provided the space is finite-dimensional and compact. Although this result represents an important milestone, the Monge problem in fully general Alexandrov spaces remains only partially understood. Beyond the Riemannian and Alexandrov settings, significant advances have been made in more general metric measure spaces. Gigli [6] obtained analogous results in synthetic Ricci curvature lower bound spaces (${\sf RCD}(K,N)$). Figalli and Rifford [5] resolved the existence and uniqueness of Monge solutions on sub-Riemannian manifolds for the squared distance cost under the assumption of absolute continuity of the source measure. The present work situates itself within this broad landscape. The proposed framework for the HJMOT problem not only encompasses the aforementioned results of Brenier, McCann, Figalli, and Bertrand, but also aligns with recent developments in general metric measure spaces. In this sense, our results provide a unified perspective that integrates and extends the classical theory of optimal transport across Euclidean, Riemannian, Alexandrov, and synthetic curvature settings.
@@
We topologize $\widehat{X}_{k}$ so that $\{\partial_{k}\}$ is isolated in $\widehat{X}_{k}$. Consequently, $\widehat{X}_{k}$ is a Polish space (see [9, Section 3]).
@@
We say that cost function $c$ is coercive if there exists a measurable function $\Phi:\Omega\to[0,+\infty]$ such that
@@
    There exists a minimizing sequence $\{\pi_{n}\}$ of (2.1) such that $\sup_{n}\int_{\Omega}\Phi\,d\pi_{n}<\infty.$
@@
If the cost function $c$ is coercive, then there exists a minimizing sequence $\{\pi_{n}\}$ of (2.1) is uniformly tight.
@@
By the coercivity condition, there exists a measurable function $\Phi:\Omega\to[0,+\infty]$ and a constant $C>0$ such that
@@
    There exists a minimizing sequence $\{\pi_{n}\}$ of (2.1) such that $\sup_{n}\int_{\Omega}\Phi\,d\pi_{n}=C.$
@@
Since $c$ is coercive, then there exists a minimizing sequence $\{\pi_{n}\}\subset\Pi(\mu_{0},\mu_{1}^{n},\cdots,\mu_{K-1}^{n},\mu_{K})$ of (2.1) with $\mu_{k}^{n}={T_{k}}_{\#}\pi_{n}$ for $k=1,\cdots,K-1$ is uniformly tight. By definition of $M$, we have
@@ 3 Existence and Uniqueness for the Monge Problem
Let $\{\pi^{*},\{\mu_{k}^{*}\}_{k=1}^{K-1}\}$ be a solution of (2.1). We say that cost function $c$ is strongly coercive if there exists a measurable function $\Phi:\Omega\to[0,+\infty]$ such that
@@
    There exists a minimizing sequence $\{\pi_{n}\}$ of (2.1) such that $\sup_{n}\int_{\Omega}\Phi\,d\pi_{n}<\infty.$
@@
    There exists $\varepsilon>0$ such that for any $\mu_{0}$-a.e. $\widehat{x}_{0}\in X_{0}$, $\{\omega\in\Omega\mid T_{0}(\omega)=\widehat{x}_{0},c(\omega)\leq h(\widehat{x}_{0})+\varepsilon\}$ is a compact set where $h(\widehat{x}_{0}):=\inf\{c(\omega)\mid\omega\in{\rm supp}(\pi^{*}),T_{0}(\omega)=\widehat{x}_{0}\}$.
Lemma 3.1 (Kantorovich duality [15]).
@@
Since $c$ is non-negative semicountinuous and $\Pi(\mu_{0}^{*},\cdots,\mu_{K}^{*})$ is uniformly tight (Lemma 2.2), then by Kellerer’s theorem ([10]) we have $D=M$ and there exists $v_{k}\in L^{1}(\widehat{X}_{k},\mu_{k}^{*})$ such that for evey $\omega\in\Omega$,
@@
Let $\{\pi^{*},\{\mu_{k}^{*}\}_{k=1}^{K-1}\}$ be a solution of (2.1), set $h(\widehat{x}_{0}):=\inf\{c(\omega)\mid\omega\in{\rm supp}(\pi^{*}),\ T_{0}(\omega)=\widehat{x}_{0}\}$. If the cost function $c$ is strongly coercive, then $S(\widehat{x}_{0}):=\{\omega\in{\rm supp}(\pi^{*})\mid T_{0}(\omega)=\widehat{x}_{0},\ c(\omega)=h(x_{0})\}$ is a non-empty set.
@@
For fix $\mu_{0}$-a.e. $\widehat{x}_{0}\in X_{0}$, by defined of $h(\widehat{x}_{0})$, there exists $\{\omega_{n}\}_{n=1}^{\infty}\subset{\rm supp}(\pi^{*})$ such that
@@
Since $c$ is strongly coercive and ${\rm supp}(\pi^{*})$ is a closed set, then there exsts $\varepsilon>0$ such that $K:=\{\omega\in{\rm supp}(\pi^{*})\mid T_{0}(\omega)=\widehat{x}_{0},c(\omega)\leq h(\widehat{x}_{0})+\varepsilon\}$ is a compact set. Then, there exists $N\in\mathbb{N}$ such that for any $n>N$,
$\left|c(\omega_{n})-h(\widehat{x}_{0})\right|<\frac{\varepsilon}{3}.$
Since $\{\omega_{n}\}_{n\geq N}^{\infty}\subset K$ and $K$ is compact, there exists $\omega_{n_{k}}$ such that $\omega_{n_{k}}\to\omega^{*}\in K$ as $k\to\infty$. For each $k$, $T_{0}(\omega_{n_{k}})=\widehat{x}_{0}$ and $T_{0}$ is continuous map, then
@@
    The coupling $\pi^{*}=T^{*}_{\#}\mu_{0}$ is the unique solution of (2.1).
@@
exists in $\mathbb{R}$. Since $c$ is locally control,
$\frac{c(\widehat{x}_{0}^{n},\widehat{x}_{1},\cdots,\widehat{x}_{K})-c(\widehat{x}_{0},\widehat{x}_{1},\cdots\,\widehat{x}_{K})}{{\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0})}=\frac{h(\widehat{x}_{0}^{n})+o({\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0}))-h(\widehat{x}_{0})}{{\sf d}_{0}(\widehat{x}_{0}^{n},\widehat{x}_{0})}.$
@@
and $D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}c(\omega)=D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}h$ for any $\omega\in S(\widehat{x}_{0})$. Suppose there exists $\omega_{1}\in S(\widehat{x}_{0})$ and $\omega_{2}\in S(\widehat{x}_{0})$ such that $\omega_{1}\neq\omega_{2}$. Since $\omega_{1}$ and $\omega_{2}$ are not equal, by $c$ is sequential twisted, $D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}c(\omega_{1})\neq D_{\{\widehat{x}_{0}^{n}\},\,\widehat{x}_{0}}c(\omega_{2})$. But $\omega_{1},\omega_{2}\in S(\widehat{x}_{0})$,
@@
contradiction. Then $S(\widehat{x}_{0})$ is a singleton. In particular, if Polish spaces $X_{0}$ that lack a nontrivial curve structure, $S(\widehat{x}_{0})=\mathcal{F}(\widehat{x}_{0})$ is also a singleton. For each $\widehat{x}_{0}\in X_{0}$, define $T^{*}(\widehat{x}_{0})$ as the unique element of $S(\widehat{x}_{0})$. Consider
$G_{r}(T^{*}):$ $=\{(\widehat{x}_{0},\omega)\in X_{0}\times\Omega\mid\omega\in S(\widehat{x}_{0})\}$
@@
Since $\{(\widehat{x}_{0},\omega)\in X_{0}\times\Omega\mid\omega\in{\rm supp}(\pi^{*}),\,T_{0}(\omega)=\widehat{x}_{0}\}$ is a closed set and $\{(\widehat{x}_{0},\omega)\in X_{0}\times\Omega\mid c(\omega)=h(\widehat{x}_{0})\}$ is a Borel set, then $G_{r}$ is a Borel set and hence $T^{*}$ is a Borel map (see [11]). By the disintegration theorem, for any Borel set $A\subset\Omega$
@@
where $\delta_{x_{0}}$ is a Dirac measure and $\kappa_{\widehat{x}_{0}}$ is the conditional probability measure of $\omega$ given $\widehat{x}_{0}$. Since $\kappa_{\widehat{x}_{0}}(S(\widehat{x}_{0}))=1$ and $S(\widehat{x}_{0})$ is a singleton, then $\kappa_{\widehat{x}_{0}}=\delta_{T^{*}(\widehat{x}_{0})}$. Therefore
$\pi^{*}(A)=\int_{X_{0}}[\delta_{\widehat{x}_{0}}\otimes\kappa_{\widehat{x}_{0}}](A)\ \mu_{0}({\rm d}\widehat{x}_{0})=T^{*}_{\#}\mu_{0}(A).$
@@
Similarly, $S(\widehat{x}_{0})$ is also a singleton and $\kappa_{\widehat{x}_{0}}(S(\widehat{x}_{0}))=1$, then $\kappa^{\prime}_{\widehat{x}_{0}}=\delta_{T^{*}(\widehat{x}_{0})}=\kappa_{\widehat{x}_{0}}$. Hence,